# A Closer Look at Time Steps is Worthy of Triple Speed-Up for Diffusion Model Training

### v1

@@ 1 Introduction
Recently, some acceleration methods for diffusion training focus on time steps, primarily using re-weighting and re-sampling 1) Re-weighting on the time steps based on heuristic rules. P2 choi2022perception and Min-SNR hang2023efficient use monotonous and single-peak weighting strategies according to sign-to-noise ratios (SNR) in different time steps. 2) Re-sampling the time steps. Log-Normal karras2022elucidating assigns high sampling probabilities for the middle time steps of the diffusion process. CLTS xu2024towards proposes a curriculum learning based time step schedule, gradually tuning the sampling probability from uniform to Gaussian by interpolation for acceleration as shown in Fig. 1(b).
@@
Previous acceleration works have achieved promising results, but the analysis of time steps remains relatively under-explored. P2 (choi2022perception, ) and Min-SNR (hang2023efficient, ) are two re-weighting methods, with their weighting curves across time steps as shown in Fig. 1(b). They employ uniform sampling of time steps, which include too many easy samples from the convergence area during diffusion model training.
On the other hand, most re-sampling methods heuristically emphasize sampling the middle-time steps, but they do not dive into the difference between the acceleration and convergence areas. For example, CLTS (xu2024towards, ) gradually changes the sampling distribution from uniform to Gaussian by interpolation as shown in Fig. 1(b). The sampling probability of the acceleration area drops faster than the one of the convergence area. The acceleration area is still under-sampled and therefore not well-learned.
Figure 1: Motivation: designing an efficient training via analyzing process increment $\delta_{t}$ at different time steps. (a) $\mathbf{E}(\delta_{t})$ and $\text{Var}(\delta_{t})$ are the mean and variance of process increments $\delta_{t}$. Two histograms represent the proportions of the process increments at different noise levels (left) and the proportions of the time steps (right) in the three areas. The loss curve is obtained from DDPM (ho2020denoising, ) on CIFAR-10 (krizhevsky2009learning, ). (b) $w(t)$ and $\mathbf{P}(t)$ are respectively the weighting and sampling curve. The probability of convergence area being sampled remains, while the one of acceleration is reduced faster.
@@
Our approach has the following characteristics: SpeeD is compatible with various diffusion model training methods, i.e., U-Net ronneberger2015u and DiT peebles2023scalable with minimal modifications. For performance, SpeeD achieves non-trivial improvements than baseline and other methods at the same training iterations. For efficiency, SpeeD consistently accelerates the diffusion training by 3$\times$ across various tasks and datasets. It helps mitigate the heavy computational cost for diffusion model training, enabling more researchers to train a model at an acceptable expense. The extra time complexity of SpeeD is $\mathcal{O}(1)$, costing only seconds to reduce days of diffusion models training on datasets like FFHQ, MetFaces and ImageNet-1K. We hope this work can bring novel insights for efficient diffusion model training.
2 Speeding Up Diffusion Training from Time Steps
@@ 2.1 Preliminaries of Diffusion Models
Our main implementation follows DDPM (ho2020denoising, ; sohl2015deep, ). Given data $x_{0}\sim p(x_{0})$, the forward process is a Markov-Gaussian process that gradually adds noise to obtain a perturbed sequence {$x_{1},x_{2},...,x_{T}$},
@@
For diffusion model training, the forward process is divided into pairs of samples and targeted process increments by time steps $t$, defined as $\delta_{t}:=x_{t+1}-x_{t}$. The diffusion model is expected to predict the next step from the given time step. The training loss ho2020denoising for diffusion models is to predict the normalized noise in a close-form. The training loss highlighted with weighting and sampling modules:
@@
2.2 A Closer Look at Time Steps
@@
In DDPM’s setting (ho2020denoising, ), the linear schedule hyper-parameters $\{\beta_{t}\}_{t\in[T]}$ is an equivariant series, the extreme deviation $\Delta_{\beta}:=\max_{t}\beta_{t}-\min_{t}\beta_{t}$, $T$ is the total number of time steps, and we have the bounds about the process increment $\delta_{t}\sim\mathcal{N}(\phi_{t},\Psi_{t})$, where $\phi_{t}:=(\sqrt{\alpha_{t+1}}-1)\sqrt{\bar{\alpha}_{t}}x_{0},\Psi_{t}:=[2-\bar{\alpha}_{t}(1+\alpha_{t+1})]\mathbf{I}$, $\mathbf{I}$ is the unit matrix, as follows:
@@
Definition of $t_{\text{d-c}}$. The process is considered to be in the convergence area where the increments’ variance is within a range. The convergence area is identified by a magnitude $r$, where $1-1/r$ is the ratio to the maximum variance. According to Theorem 1, the convergence area is defined as one magnitude reduction of the scale factor (i.e., $1\times r$), and we have the lower-bound of the magnitude $\hat{r}_{t}:=\exp\{(\beta_{0}+\Delta_{\beta}t/2T)t\}$ employed as the threshold selection function in Section 2.4. The time step $t$ is guaranteed to be in convergence area when $\hat{r}_{t}>r$, .
@@
In the convergence area, the variations of $\delta_{t}$ stabilize, signifying that the process is approaching a steady state. This area corresponds to a very large proportion of the overall time steps. On top of that, the training loss in this area is empirically low, which leads to the redundant training cost on time steps with limited benefits for training. In the acceleration area, the variations of $\delta_{t}$ increase, indicating a rapid change. Conversely, in the deceleration area, the variations of $\delta_{t}$ decrease, reflecting a slowing down of the process. Notably, near the peak between the acceleration and deceleration areas, the process exhibits the fastest changes. These time steps only occupy a small proportion. Beyond that, the training losses in this area are empirically high. The issue is that a hard-to-learn area is even under-sampled, necessitating more sampling and training efforts.
@@ 3 Experiments
We first describe the mian experiment setups in Sec. 3.1, including benchmark datasets, network architectures, training details, and evaluation protocols. In Sec. 3.2, we present the main results regarding both performance and efficiency. After that, in Sec. 3.5 and 3.3, we ablate the effectiveness of each designed module and evaluate the generalization of SpeeD in various diffusion tasks and settings. Finally, we investigate the compatibility of our approach with other recent methods in 3.4 and present visualizations in Sec. 3.6.
@@ 3.1 Implementation Details
Datasets. We mainly investigate the effectiveness of our approach on the following datasets: MetFaces (karras2020training, ) and FFHQ (karras2019style, ) are used to train unconditional image generation, CIFAR-10 (krizhevsky2009learning, ) and ImageNet-1K (deng2009imagenet, ) are used to train conditional image generation, and MS-COCO (lin2014microsoft, ) is used to evaluate the generalization of our method in the text to image task. More details of these datasets can be found in the Appendix A.
Network architectures. U-Net ronneberger2015u and DiT peebles2023scalable are two famous architectures in the diffusion model area. We implement our approach on these two architectures and their variants. We follow the same hyper parameters as the baseline by default. More information about the details of the architectures can be found in Appendix A.1.
Training details. We train all models using AdamW  kingma2014adam ; loshchilov2017decoupled with a constant learning rate 1e-4. We set the maximum step in training to 1000 and use the linear variance. All images are augmented with horizontal flip transformations if not stated otherwise. Following common practice in the generative modeling literature, the exponential moving average (EMA) (gardner1985exponential, ) of network weights is used with a decay of 0.9999. The results are reported using the EMA model. Details can be found in Tab. 6.
@@
dataset MetFaces FFHQ iterations 10K 20K 30K 40K 50K 10K 20K 30K 40K 50K DiT-XL/2 peebles2023scalable 398.7 132.7 74.7 36.7 29.3 356.1 335.3 165.2 35.8 12.9 P2 choi2022perception 377.1 328.2 111.0 27.3 23.3 368.7 357.9 158.6 35.5 15.0 Min-SNR hang2023efficient 389.4 313.9 52.1 31.3 28.6 376.3 334.1 151.2 34.0 12.6 Log-Normal esser2024scaling 311.8 165.1 63.9 51.1 47.3 307.6 165.1 63.9 51.1 47.3 CLTS xu2024towards 375.0 57.2 28.6 24.6 23.5 336.1 329.1 173.4 33.7 12.7 SpeeD 367.3 23.4 22.6 22.1 21.1 322.1 320.0 91.8 19.8 9.9
@@
Before our comparison, we first introduce our baseline, i.e., DiT-XL/2, a strong image generation backbone as introduced in DiT peebles2023scalable . We follow the hyperparameter settings from DiT and train DiT-XL/2 on MetFaces karras2020training and FFHQ karras2019style , respectively. We compare our approach with two re-weighting methods: P2 choi2022perception and Min-SNR hang2023efficient , and two re-sampling methods: Log-Normal karras2022elucidating and CLTS xu2024towards . In the evaluation, we use 10K generated images to calculate FID heusel2017gans for comparison. To make a detailed comparison, all the approaches are trained for 50K iterations and we report the FID scores per 10K iterations.
@@
In addition to the performance comparison, we also present the acceleration results of our SpeeD . This naturally raises a question: how to calculate the acceleration? Here, we follow the previous diffusion acceleration methods gao2023masked and other efficient training papers qin2023infobatch ; xu2024towards : visualizing the FID-Iteration curve and reporting the estimated highest acceleration ratio. We mainly compare with DiT-XL/2, one re-weighting method Min-SNR choi2022perception , and one re-sampling method CLTS xu2024towards in Fig 4. At the same training iterations, our approach achieves significantly better FID scores than other methods. Notably, SpeeD accelerates the Min-SNR, and CLTS by 2.7 and 2.6 times, respectively. More comparisons with other methods can be found in Appendix B.
@@ 3.3 Generalization Evaluation
Cross-architecture robustness evaluation. There are mainly two architectures in the diffusion models: U-Net ronneberger2015u and DiT peebles2023scalable . SpeeD is not correlated to specific model architecture, thereby it is a model-agnostic approach. We implement our method with DiT-XL/2 and U-Net on MetFaces, FFHQ, and ImageNet-1K, respectively. We default to training the models for 50K iterations on MetFaces and FFHQ, 400K on ImageNet-1K. To ensure a fair comparison, we keep all hyper-parameters the same and report the FID scores at 50K iterations. As shown in Tab. 3, SpeeD consistently achieve significantly higher performance under all settings, which indicates the strong generality of our method for different architectures and datasets.
Cross-schedule robustness evaluation. In the diffusion process, there are various time step schedules, including linear (ho2020denoising, ), quadratic and cosine (nichol2021improved, ) schedules. We verify SpeeD’s robustness across these schedules. We report FID and inception score (IS) (salimans2016improved, ) scores as metrics for comparisons. As shown in Tab. 3, SpeeD achieves significant improvement on linear, quadratic, and cosine schedules both in FID and IS. That shows the strong generality of SpeeD in various schedules.
@@
Cross-task robustness evaluation. We apply SpeeD to the text-to-image task for evaluating the generality of our method. For text-to-image generation, we first introduce CLIP (radford2021learning, ) to extract the text embedding for MS-COCO (lin2014microsoft, ) dataset. Then, DiT-XL/2 is used to train a text-to-image model as our baseline. Following prior work(saharia2022photorealistic, ), FID score and CLIP score on MS-COCO validation set are evaluation metrics for quantitative analyses. As illustrated in Tab. 4, we obtain the better FID and CLIP score than our baseline.

### v2

@@ 1 Introduction
Recently, some acceleration methods for diffusion training focus on time steps, primarily using re-weighting and re-sampling 1) Re-weighting on the time steps based on heuristic rules. P2 (Choi et al., 2022) and Min-SNR (Hang et al., 2023) use monotonous and single-peak weighting strategies according to sign-to-noise ratios (SNR) in different time steps. 2) Re-sampling the time steps. Log-Normal (Karras et al., 2022) assigns high sampling probabilities for the middle time steps of the diffusion process. CLTS (Xu et al., 2024) proposes a curriculum learning based time step schedule, gradually tuning the sampling probability from uniform to Gaussian by interpolation for acceleration as shown in Fig. 1(b).
@@
Previous acceleration works have achieved promising results, but the analysis of time steps remains relatively under-explored. P2 (Choi et al., 2022) and Min-SNR (Hang et al., 2023) are two re-weighting methods, with their weighting curves across time steps as shown in Fig. 1(b). They employ uniform sampling of time steps, which include too many easy samples from the convergence area during diffusion model training.
On the other hand, most re-sampling methods heuristically emphasize sampling the middle-time steps, but they do not dive into the difference between the acceleration and convergence areas. For example, CLTS (Xu et al., 2024) gradually changes the sampling distribution from uniform to Gaussian by interpolation as shown in Fig. 1(b). The sampling probability of the acceleration area drops faster than the one of the convergence area. The acceleration area is still under-sampled and therefore not well-learned.
Figure 1: Motivation: designing an efficient training via analyzing process increment $\delta_{t}$ at different time steps. (a) $\mathbf{E}(\delta_{t})$ and $\text{Var}(\delta_{t})$ are the mean and variance of process increments $\delta_{t}$. Two histograms represent the proportions of the process increments at different noise levels (left) and the proportions of the time steps (right) in the three areas. The loss curve is obtained from DDPM (Ho et al., 2020) on CIFAR-10 (Krizhevsky et al., 2009). (b) $w(t)$ and $\mathbf{P}(t)$ are respectively the weighting and sampling curve. The probability of convergence area being sampled remains, while the one of acceleration is reduced faster.
@@
Our approach has the following characteristics: SpeeD is compatible with various diffusion model training methods, i.e., U-Net (Ronneberger et al., 2015) and DiT (Peebles & Xie, 2023) with minimal modifications. For performance, SpeeD achieves non-trivial improvements than baseline and other methods at the same training iterations. For efficiency, SpeeD consistently accelerates the diffusion training by 3$\times$ across various tasks and datasets. It helps mitigate the heavy computational cost for diffusion model training, enabling more researchers to train a model at an acceptable expense. The extra time complexity of SpeeD is $\mathcal{O}(1)$, costing only seconds to reduce days of diffusion models training on datasets like FFHQ, MetFaces and ImageNet-1K. We hope this work can bring novel insights for efficient diffusion model training.
2 Speeding Up Training from Process Increments at Time Steps
@@ 2.1 Preliminaries of Diffusion Models
In conventional DDPM (Ho et al., 2020; Sohl-Dickstein et al., 2015), given data $x_{0}\sim p(x_{0})$, the forward process is a Markov-Gaussian process that gradually adds noise to obtain a perturbed sequence {$x_{1},x_{2},...,x_{T}$},
@@
For diffusion model training, the forward process is divided into pairs of samples and targeted process increments by time steps $t$, defined as $\delta_{t}:=x_{t+1}-x_{t}$. The diffusion model is expected to predict the next step from the given time step. The training loss (Ho et al., 2020) for diffusion models is to predict the normalized noise. The loss highlighted with weighting and sampling modules:
@@
2.2 Case Study on DDPM: A Closer Look at Time Steps
@@
In DDPM’s setting (Ho et al., 2020), the linear schedule hyper-parameters $\{\beta_{t}\}_{t\in[T]}$ is an equivariant series, the extreme deviation $\Delta_{\beta}:=\max_{t}\beta_{t}-\min_{t}\beta_{t}$, $T$ is the total number of time steps, and we have the bounds about the process increment $\delta_{t}\sim\mathcal{N}(\phi_{t},\Psi_{t})$, where $\phi_{t}:=(\sqrt{\alpha_{t+1}}-1)\sqrt{\bar{\alpha}_{t}}x_{0},\Psi_{t}:=[2-\bar{\alpha}_{t}(1+\alpha_{t+1})]\mathbf{I}$, $\mathbf{I}$ is the unit matrix:
@@
Definition of $t_{\text{d-c}}$. The process is considered to be in the convergence area where the increments’ variance is within a range. The convergence area is identified by a magnitude $r$, where $1-1/r$ is the ratio to the maximum variance.
According to Theorem 1, the convergence area is defined as one magnitude reduction of the scale factor (i.e., $1\times r$), and we have the lower-bound of the magnitude $\hat{r}_{t}:=\exp\{(\beta_{0}+\Delta_{\beta}t/2T)t\}$ employed as the threshold selection function in Section 2.5. The time step $t$ is guaranteed to be in convergence area when $\hat{r}_{t}>r$, .
@@
In the convergence area, the variations of $\delta_{t}$ stabilize, signifying that the process is approaching a steady state.
This area corresponds to a very large proportion of the overall time steps. On top of that, the training loss in this area is empirically low, which leads to the redundant training cost on time steps with limited benefits for training.
In the acceleration area, the variations of $\delta_{t}$ increase, indicating a rapid change. Conversely, in the deceleration area, the variations of $\delta_{t}$ decrease, reflecting a slowing down of the process. Notably, near the peak between the acceleration and deceleration areas, the process exhibits the fastest changes. These time steps only occupy a small proportion. Beyond that, the training losses in this area are empirically high. The issue is that a hard-to-learn area is even under-sampled, necessitating more sampling and training efforts.
@@
2.3 General Cases Beyond DDPM: Dive into The Commonality.
This section generalize the above findings on DDPM to broader settings. The findings are about the process increments $\delta_{t}:=x_{t+1}-x_{t}$, and the related differentiation of right limit $\mathbf{d}x$ in forward process. The corresponding SDE (Karras et al., 2022) and the discretization are:
$\mathbf{d}x=x\dot{s}/s\mathbf{d}t+s\sqrt{\dot{\sigma}\sigma}\mathbf{d}w,\qquad x_{t}=s_{t}x_{0}+s_{t}\sigma_{t}\epsilon,\epsilon\sim\mathcal{N}(0,\mathbf{I}),$
where scale factor $s=s_{t}$ and noise standard deviation (std.) $\sigma=\sigma_{t}$ are the main designs related to the main factors $\hat{\phi}_{t}$ and $\hat{\Psi}_{t}$ about process increment $\delta_{t}$ in Theorem 1 across time steps $t$.
Generalize Theorem 1: $s$-$\sigma$ Scheduled Process Increments.
The generalized process increment is $\delta\sim\mathcal{N}(\Delta x_{0},\Sigma\mathbf{I})$, where $\Delta:=s_{+}-s$ and $\Sigma:=s_{+}^{2}\sigma_{+}^{2}+s^{2}\sigma^{2}$ across $t$. $\Delta$, $\Sigma$ are continuous on $t$ without discretization, where $s_{+}$ and $\sigma_{+}$ are the right outer limits, i.e., $s(t+\mathbf{d}t)$. In discretization, $\Delta_{t}$ and $\Sigma_{t}$, marked by $t$, are related to sample granularity of time step $t$. Like Theorem 1, we study the variation of process increments by $\dot{\Delta}=\dot{s}_{+}-\dot{s}$ and $\dot{\Sigma}=\mathbf{m}^{\top}\dot{\mathbf{n}}$ where $\mathbf{m}=[s_{+}^{2},\sigma_{+}^{2},s^{2},\sigma^{2}]^{\top}$, $\mathbf{n}=[\sigma_{+}^{2},s_{+}^{2},\sigma^{2},s^{2}]^{\top}$. This formulation involves only terms about derivatives of given schedule functions, which brings computational convenience. Tab. 1 provides all ingredients needed to calculate curves of $\dot{\Delta}$ and $\dot{\Sigma}$ in schedules of VP, VE and EDM.
We also generalize the previous takeaways from the DDPM case study to $s$-$\sigma$ scheduled setting with the following analyses and discussion.
$\sigma$ is better for the design of sampling and weighting than $s$. It stands due to its direct reflection about SNR (signal-to-noise ratio), and additionally, because $s$ is usually adapted to heuristic motivations. In DDPM, corresponding to VP in Tab. 1, the SDE design is simply from data to normal noise. In VE, realistic diffusion processes inspire that the diffusion rate is limited to $\sigma=\sqrt{t}$. Further in EDM, motivation become more complex of training objective and concise of schedule and motivation at the same time, bringing benefits to training. Its key ideas and designs are 1) the std. of inputs and targeted outputs of a neural network $F_{\theta}$ in EDM is constrained to 1 with preconditioning; 2) the weights $w_{t}$ in Eqn. 1 are allocated according to $c_{\text{out}}w_{t}=1$, where $c_{\text{out}}$ is the scale factor of F-prediction neural network $F_{\theta}$’s outputs, and is related to $\sigma$ and the std. of data (sometimes normalized as 1).
Sampling deserves more attention. The SDE design goal of most diffusion models nowadays is to add much larger noise to the data so that the samples can cover larger space. However, in terms of the process increment, it always results in a low signal-to-noise ratio of the late data when $t$ is large. Either the standard deviation is too large or the $s$ used to suppress it is too small. For instance, EDM does not bias the data distribution from expectations due to the scale $\dot{\Delta}=0$, but $\dot{\Sigma}=2[(t+\mathbf{d}t)^{2}+t^{2}]$ is a quadratic increase as $t$ grows. In VP, $s\rightarrow 0$, as $t$ grows, leads that the model needs to recover expectations in approximation. For these samples, which are not very informative, a single weight adjustment is not as efficient as directly reducing the sampling rate.
Table 1: The ingredients of generalized curves $\dot{\Delta}$ and $\dot{\Sigma}$ schedules about mainstream SDE designs, including VP, VE (Song et al., 2021), EDM (Karras et al., 2022).
Schedules $s$ $\sigma^{2}$ $\dot{s}$ $\dot{\sigma}$
VP $\exp\{-\frac{1}{4}\Delta_{\beta}t^{2}-\frac{1}{2}\beta_{0}t\}$ $\exp\{\frac{1}{2}\Delta_{\beta}t^{2}+\beta_{0}t\}-1$ $-\frac{\sigma\dot{\sigma}}{(1+\sigma^{2})^{3/2}}$ $\frac{(1+\sigma^{2})(\Delta_{\beta}t+\beta_{0})}{2\sigma}$
VE $1$ $t$ 00 $1$
EDM $1$ $t^{2}$ 00 $2t$
@@ 3 Experiments
We first present visualizations in Sec. 3.1 and describe the main experiment setups in Sec. 3.2, including benchmark datasets, network architectures, training details, and evaluation protocols. In Sec. 3.3, we present the main results regarding both performance and efficiency. After that, in Sec. 3.6 and 3.4, we ablate the effectiveness of each designed module and evaluate the generalization of SpeeD in various diffusion tasks and settings. Finally, we investigate the compatibility of our approach with other recent methods in 3.5.
3.1 Visualization
(a) Visualization comparison on MetFaces dataset.
(b) Visualization comparison on FFHQ dataset.
Figure 4: Our SpeeD obtains significant improvements than baseline in visualizations. More visualizations on other datasets and tasks can be found in the Appendix D.
@@ 3.2 Implementation Details
Datasets. We mainly investigate the effectiveness of our approach on the following datasets: MetFaces (Karras et al., 2020) and FFHQ (Karras et al., 2019) are used to train unconditional image generation, CIFAR-10 (Krizhevsky et al., 2009) and ImageNet-1K (Deng et al., 2009) are used to train conditional image generation, and MS-COCO (Lin et al., 2014) is used to evaluate the generalization of our method in the text to image task. More details of these datasets can be found in the Appendix A.
Network architectures. U-Net (Ronneberger et al., 2015) and DiT (Peebles & Xie, 2023) are two famous architectures in the diffusion model area. We implement our approach on these two architectures and their variants. We follow the same hyper parameters as the baseline by default. More information about the details of the architectures can be found in Appendix A.3.
Training details. We train all models using AdamW  (Kingma & Ba, 2014; Loshchilov & Hutter, 2017) with a constant learning rate 1e-4. We set the maximum step in training to 1000 and use the linear variance. All images are augmented with horizontal flip transformations if not stated otherwise. Following common practice in the generative modeling literature, the exponential moving average (EMA) (Gardner Jr, 1985) of network weights is used with a decay of 0.9999. The results are reported using the EMA model. Details can be found in Tab. 7.
@@
dataset MetFaces FFHQ iterations 10K 20K 30K 40K 50K 10K 20K 30K 40K 50K DiT-XL/2 (Peebles & Xie, 2023) 398.7 132.7 74.7 36.7 29.3 356.1 335.3 165.2 35.8 12.9 P2 (Choi et al., 2022) 377.1 328.2 111.0 27.3 23.3 368.7 357.9 158.6 35.5 15.0 Min-SNR (Hang et al., 2023) 389.4 313.9 52.1 31.3 28.6 376.3 334.1 151.2 34.0 12.6 Log-Normal (Esser et al., 2024) 311.8 165.1 63.9 51.1 47.3 307.6 165.1 63.9 51.1 47.3 CLTS (Xu et al., 2024) 375.0 57.2 28.6 24.6 23.5 336.1 329.1 173.4 33.7 12.7 SpeeD (ours) 367.3 23.4 22.6 22.1 21.1 322.1 320.0 91.8 19.8 9.9
@@
Before our comparison, we first introduce our baseline, i.e., DiT-XL/2, a strong image generation backbone as introduced in DiT (Peebles & Xie, 2023). We follow the hyperparameter settings from DiT and train DiT-XL/2 on MetFaces (Karras et al., 2020) and FFHQ (Karras et al., 2019), respectively. We compare our approach with two re-weighting methods: P2 (Choi et al., 2022) and Min-SNR (Hang et al., 2023), and two re-sampling methods: Log-Normal (Karras et al., 2022) and CLTS (Xu et al., 2024). In the evaluation, we use 10K generated images to calculate FID (Heusel et al., 2017) for comparison. To make a detailed comparison, all the approaches are trained for 50K iterations and we report the FID scores per 10K iterations.
@@
In addition to the performance comparison, we also present the acceleration results of our SpeeD . This naturally raises a question: how to calculate the acceleration? Here, we follow the previous diffusion acceleration methods (Gao et al., 2023) and other efficient training papers (Qin et al., 2023; Xu et al., 2024): visualizing the FID-Iteration curve and reporting the estimated highest acceleration ratio. We mainly compare with DiT-XL/2, one re-weighting method Min-SNR (Choi et al., 2022), and one re-sampling method CLTS (Xu et al., 2024) in Fig 5. At the same training iterations, our approach achieves significantly better FID scores than other methods. Notably, SpeeD accelerates the Min-SNR, and CLTS by 2.7 and 2.6 times, respectively. More comparisons with other methods can be found in Appendix B.
@@ 3.4 Generalization Evaluation
Cross-architecture robustness evaluation. There are mainly two architectures in the diffusion models: U-Net (Ronneberger et al., 2015) and DiT (Peebles & Xie, 2023). SpeeD is not correlated to specific model architecture, thereby it is a model-agnostic approach. We implement our method with DiT-XL/2 and U-Net on MetFaces, FFHQ, and ImageNet-1K, respectively. We default to training the models for 50K iterations on MetFaces and FFHQ, 400K on ImageNet-1K.To ensure a fair comparison, we keep all hyper-parameters the same and report the FID scores at 50K iterations. As shown in Tab. 4, SpeeD consistently achieve significantly higher performance under all settings, which indicates the strong generality of our method for different architectures and datasets.
Cross-schedule robustness evaluation. In the diffusion process, there are various time step schedules, including linear (Ho et al., 2020), quadratic and cosine (Nichol & Dhariwal, 2021) schedules. We verify SpeeD’s robustness across these schedules. We report FID and inception score (IS) (Salimans et al., 2016) scores as metrics for comparisons. As shown in Tab. 4, SpeeD achieves significant improvement on linear, quadratic, and cosine schedules both in FID and IS. That shows the strong generality of SpeeD in various schedules.
@@
Cross-task robustness evaluation. We apply SpeeD to the text-to-image task for evaluating the generality of our method. For text-to-image generation, we first introduce CLIP (Radford et al., 2021) to extract the text embedding for MS-COCO (Lin et al., 2014) dataset. Then, DiT-XL/2 is used to train a text-to-image model as our baseline. Following prior work(Saharia et al., 2022), FID score and CLIP score on MS-COCO validation set are evaluation metrics for quantitative analyses. As illustrated in Tab. 5, we obtain the better FID and CLIP score than our baseline.